# IVE: Enhanced Probabilistic Forecasting of Intraday Volume Ratio with Transformers

### v1

@@ 1 Introduction
In the dynamic realm of financial markets, precise execution strategies, notably the Volume-Weighted Average Price (VWAP), are essential Almgren and Chriss (2001); Kato (2015). Success in VWAP hinges on accurate volume ratio predictions, a complex challenge given market unpredictability. Addressing this, our research presents an new approach, leveraging the Transformer architecture Vaswani et al. (2023) for high accuracy in intraday volume ratio predictions at a minute-level granularity.
Diverging from traditional models Ma and Li (2021); Liu and Lai (2016) reliant on log-transformed volumes or turnover rates, our model incorporates a log-normal transformation to counter the high variability of intraday volume ratios. We enrich our model with a diverse dataset, including external volume-related features, temporal information, and stock-specific characteristics, enhancing predictive accuracy.
@@ 2.1 Optimal Execution and VWAP
A key performance measure in optimal execution strategies is the Volume Weighted Average Price (VWAP). VWAP, influenced by the actions of all market participants, measures the average transaction price of all trades over a time period. The challenge in targeting VWAP stems from its dependency on the uncertain actions of other market participants and the inherent randomness in future market order-flows. Kato (2015)
An emerging focus in optimal execution, especially within VWAP strategies, is the importance of precise, short-term predictions. The ability to accurately forecast market conditions on smaller time scales, such as minute-by-minute changes, is crucial for effective implementation of these strategies. This precision allows for more efficient order execution, better aligning with the constantly fluctuating market conditions. Dang and Chen (2013)
Various VWAP strategies have been developed to follow this benchmark, emphasizing mathematical modeling of factors like the percentage of Volume (POV), percentage of cumulative volume (POCV), and turnover rate. These models aim to distribute the target quantity appropriately at each time point, considering the short-term market dynamics. Kato (2015)
In VWAP-conforming optimal execution strategies, breaking down large orders into smaller ones for execution over shorter periods at better than average market prices is a key tactic. This method is particularly beneficial for large-scale orders, as it reduces market impact and uses price volatility to minimize transaction costs. Kato (2015)
@@ 2.2 Intraday Volume Prediction
The U-shaped intraday volume pattern has long been observed, and statistical modeling has historically been used to predict this U-shape and the volume curve. Karpoff (1986) Past studies have used statistical modeling techniques (e.g., ARIMA) and traditional machine learning approaches for prediction at different scales, including 50-minute, 10-minute, and up to 5-minute intervals. Graczyk and Duarte Queirós (2016)
These studies have indicated that gradient-based methodologies can be effective for precise predictions, and RNN-based approaches have been successfully implemented on a 50-minute scale. However, RNNs have limitations due to gradient vanishing, making them less suitable for shorter scales. Libman et al. (2019)
In this paper, we effectively perform minute-scale probabilistic intraday volume forecasting using a Transformer-based structure. This approach allows for the construction of simple volume distribution strategies and superior performance in VWAP target strategies. Graczyk and Duarte Queirós (2016)
@@ 3 Methodology and Experiments
3.1 Data Creation and Feature Engineering
@@
Delving deeper into the nuances of trading volume analysis, it becomes apparent that the simple statistical characteristics of volume alone are insufficient to capture the complexity of market behaviors. Graczyk and Duarte Queirós (2016) Our exploration reveals that trading volume curves are significantly shaped by a myriad of factors, including intraday seasonalities, nonstationarities Graczyk and Duarte Queirós (2016), and absolute time information Libman et al. (2019). Furthermore, the liquidity levels of stocks introduce another layer of complexity, as evidenced by the disparate volume curves observed in stocks with varying liquidity Markov et al. (2019). The consistent pattern of trading volume curves, despite these variations, emphasizes the importance of considering historical trading data, as evidenced in previous studies. Liu and Lai (2016); Markov et al. (2019)
This understanding guided the development of our model, named IVE (Intraday Volume Estimator), with its input features tailored to maximize the efficiency of the Transformer Architecture. By integrating Time Encoding alongside the standard sinusoidal positional embedding, the model gains enhanced capabilities in processing absolute time information. The decision to set the context length at 390, encompassing a full day’s historical data, stems from the intention to provide a comprehensive snapshot of trading activities. This includes an array of volume-related features such as volume, accumulated volume, turnover rate, and trade amount, enabling a more nuanced understanding of the market dynamics.
@@
3.2 Model Achitecture and Training
In this section, we delve into the model architecture and training methodology for predicting trading volume ratios using a transformer-based approach. Inspired by the Hugging Face blog Rogge and Rasul (2022), our model incorporates distribution heads into the transformer architecture to probabilistically predict trading volume ratios.
Input encoding methods for deep-learning approach for time series modeling usually encompass two main approaches: linear layers and convolutional layers. While 1D convolutional layers have been prevalent when modeling financial data with transformers Wallbridge (2020), they are more effective for low-dimensional multivariate data. Given our focus on high-dimensional multivariate data, we opted for linear projection to encode input features. Zerveas et al. (2020)
The transformed input features are subsequently processed by a transformer architecture. However, instead of a standard output layer, our model employs a distribution head that models the output as a T-distribution. This choice is motivated by the observation from quantitative research  Graczyk and Duarte Queirós (2016) that trading volume follows a fat-tailed stationary distribution in financial markets, making the T-distribution a more suitable choice compared to a normal distribution.
To prevent trivial solutions during training, we set our objective to predict not only one step ahead but up to three steps ahead. This approach addresses the tendency of time series transformers to impute missing values by predicting based on previous values. Zerveas et al. (2020) Nevertheless, in practical applications, we utilize predictions up to one step ahead for trading decisions.
@@ 3.2 Model Achitecture and Training
To assess the effectiveness of our approach, we compare it with several baseline algorithms commonly used in previous research. These include bi-directional LSTM (Bi-LSTM), LSTM,and recurrent neural network (RNN). Previous studies often relied on hidden state models like Kalman filters, statistical models like ARMA and GARCH, and machine learning techniques like Gradient Boosting. Satish et al. (2014); Ma and Li (2021); Antulov-Fantulin et al. (2020); Liu and Lai (2016) However, these methods were tailored to the statistical properties of trading volume curves, which cannot provide the context needed for predicting trading volume ratios.
LSTM-HR-AR, introduced in  (Zhang et al., 2019), provided a notable precedent for using deep learning to predict trading volume ratios. However, the AR model in their architecture cannot be applied to our task of predicting trading volume ratios, leading us to select the LSTM-HR model for comparison. In the interest of a fair comparison, we constructed RNN and Bi-LSTM models similarly to the LSTM-HR model.
@@
Table 1: Experimental results for the Korean market.
@@
Table 2: Experimental results for the U.S. market.
@@ 4 Possibilities of Intraday Volume Spike Detection
Integrating these predictions with other indicators could enhance the practical application of volume ratio prediction-based order execution algorithms. In our real-world trading tests, we conducted an analysis using historical data covering eight months to extract the median of volume ratio increases. We then developed a logic utilizing the model’s standard deviation (std) specifically for cases exceeding this median. Notably, in these live trading tests, we observed that the model demonstrated predictive power where the increase in volume ratio was higher than the median. This observation offers insights into the practical utility of model’s uncertainty towards volume spike in actual market scenarios. To account for sudden jumps in predicted values, we incorporated a constant factor of approximately 0.2, which was multiplied by the standard deviation and added to the predicted values. This constant was empirically adjusted based on observed performance.
@@ 6 Conclusion
In conclusion, our study presents a advancement in the field of intraday volume ratio prediction, integrating transformer architecture with practical trading applications.
7 ACKNOWLEDGEMENT
@@ A.1.1 Data and Features
We utilized execution performance data from September 26 to November 21. The features used in the analysis are as follows:
@@ A.1.2 Multiple Regression Analysis Results
The results of the multiple regression analysis are summarized in the table below:
@@ A.1.3 Interpretation of Results
  • min_bar_low2high_std (x4): The algorithm performs better in conditions of high volatility within minute bars This implies that IVE is effective in environments with frequent extreme price movements.

### v2

@@ 1 Introduction
In the dynamic realm of financial markets, precise execution strategies, notably the Volume-Weighted Average Price (VWAP), are essential Almgren & Chriss (2001); Kato (2015). Success in VWAP hinges on accurate volume ratio predictions, a complex challenge given market unpredictability. Addressing this, our research presents a new approach, leveraging the Transformer architecture Vaswani et al. (2023) for high accuracy in intraday volume ratio predictions at a minute-level granularity.
Diverging from traditional models Ma & Li (2021); Liu & Lai (2016) reliant on log-transformed volumes or turnover rates, our model incorporates a log-normal transformation to counter the high variability of intraday volume ratios. We enrich our model with a diverse dataset, including external volume-related features, temporal information, and stock-specific characteristics, enhancing predictive accuracy.
@@ 2.1 Optimal Execution and VWAP
A key performance measure in optimal execution strategies is the Volume Weighted Average Price (VWAP). VWAP, influenced by the actions of all market participants, measures the average transaction price of all trades over a time period. The challenge in targeting VWAP stems from its dependency on the uncertain actions of other market participants and the inherent randomness in future market order-flows Kato (2015).
An emerging focus in optimal execution, especially within VWAP strategies, is the importance of precise, short-term predictions. The ability to accurately forecast market conditions on smaller time scales, such as minute-by-minute changes, is crucial for effective implementation of these strategies. This precision allows for more efficient order execution, better aligning with the constantly fluctuating market conditions Dang & Chen (2013).
Various VWAP strategies have been developed to follow this benchmark, emphasizing mathematical modeling of factors like the percentage of Volume (POV), percentage of cumulative volume (POCV), and turnover rate. These models aim to distribute the target quantity appropriately at each time point, considering the short-term market dynamics Kato (2015).
In VWAP-conforming optimal execution strategies, breaking down large orders into smaller ones for execution over shorter periods at better than average market prices is a key tactic. This method is particularly beneficial for large-scale orders, as it reduces market impact and uses price volatility to minimize transaction costs Kato (2015).
@@ 2.2 Intraday Volume Prediction
The U-shaped intraday volume pattern has long been observed, and statistical modeling has historically been used to predict this U-shape and the volume curve Karpoff (1986). Past studies have used statistical modeling techniques (e.g., ARIMA) and traditional machine learning approaches for prediction at different scales, including 50-minute, 10-minute, and up to 5-minute intervals Graczyk & Duarte Queirós (2016).
These studies have indicated that gradient-based methodologies can be effective for precise predictions, and RNN-based approaches have been successfully implemented on a 50-minute scale. However, RNNs have limitations due to gradient vanishing, making them less suitable for shorter scales Libman et al. (2019).
In this paper, we effectively perform minute-scale probabilistic intraday volume forecasting using a Transformer-based structure. This approach allows for the construction of simple volume distribution strategies and superior performance in VWAP target strategies Graczyk & Duarte Queirós (2016).
@@ 3 Methodology and Experiments
3.1 Data Creation
@@
Delving deeper into the nuances of trading volume analysis, it becomes apparent that the simple statistical characteristics of volume alone are insufficient to capture the complexity of market behaviors Graczyk & Duarte Queirós (2016). Our exploration reveals that trading volume curves are significantly shaped by a myriad of factors, including intraday seasonalities, nonstationarities Graczyk & Duarte Queirós (2016), and absolute time information Libman et al. (2019). Furthermore, the liquidity levels of stocks introduce another layer of complexity, as evidenced by the disparate volume curves observed in stocks with varying liquidity Markov et al. (2019). The consistent pattern of trading volume curves, despite these variations, emphasizes the importance of considering historical trading data, as evidenced in previous studies Liu & Lai (2016); Markov et al. (2019).
This understanding guided the development of our model, named IVE (Intraday Volume Estimator), with its input features tailored to maximize the efficiency of the Transformer Architecture. By integrating Time Encoding alongside the standard sinusoidal positional embedding, the model gains enhanced capabilities in processing absolute time information. The decision to set the context length at 390 one-minute steps, encompassing a full day’s historical data, stems from the intention to provide a comprehensive snapshot of trading activities. This includes an array of volume-related features such as volume, accumulated volume, turnover rate, and trade amount, enabling a more nuanced understanding of the market dynamics.
@@
3.2 Model Architecture and Training
In this section, we delve into the model architecture and training methodology for predicting trading volume ratios using a transformer-based approach. Inspired by the Hugging Face blog Rogge & Rasul (2022), our model incorporates distribution heads into the transformer architecture to probabilistically predict trading volume ratios.
Input encoding methods for deep-learning approach for time series modeling usually encompass two main approaches: linear layers and convolutional layers. While 1D convolutional layers have been prevalent when modeling financial data with transformers Wallbridge (2020), they are more effective for low-dimensional multivariate data. Given our focus on high-dimensional multivariate data, we opted for linear projection to encode input features Zerveas et al. (2020).
The transformed input features are subsequently processed by a transformer architecture. However, instead of a standard output layer, our model employs a distribution head that models the output as a Student’s t-distribution. This choice is motivated by the observation from quantitative research Graczyk & Duarte Queirós (2016) that trading volume follows a fat-tailed stationary distribution in financial markets, making the t-distribution a more suitable choice compared to a normal distribution.
To prevent trivial solutions during training, we set our objective to predict not only one step ahead but up to three steps ahead. This approach addresses the tendency of time series transformers to impute missing values by predicting based on previous values Zerveas et al. (2020). Nevertheless, in practical applications, we utilize predictions up to one step ahead for trading decisions.
@@ 3.2 Model Architecture and Training
To assess the effectiveness of our approach, we compare it with several baseline algorithms commonly used in previous research. These include bi-directional LSTM (Bi-LSTM), LSTM, and recurrent neural network (RNN). Previous studies often relied on hidden state models like Kalman filters, statistical models like ARMA and GARCH, and machine learning techniques like Gradient Boosting Satish et al. (2014); Ma & Li (2021); Antulov-Fantulin et al. (2020); Liu & Lai (2016). However, these methods were tailored to the statistical properties of trading volume curves, which cannot provide the context needed for predicting trading volume ratios.
LSTM-HR-AR, introduced in Zhang et al. (2019), provided a notable precedent for using deep learning to predict trading volume ratios. However, the AR model in their architecture cannot be applied to our task of predicting trading volume ratios, leading us to select the LSTM-HR model for comparison. In the interest of a fair comparison, we constructed RNN and Bi-LSTM models similarly to the LSTM-HR model.
@@
Table 1: Experimental results for the Korean market (MAE in VWAP units).
@@
Table 2: Experimental results for the U.S. market (MAE in VWAP units).
@@ 4 Possibilities of Intraday Volume Spike Detection
Integrating these predictions with other indicators could enhance the practical application of volume ratio prediction-based order execution algorithms. In our real-world trading tests, we conducted an analysis using historical data covering eight months to extract the median of volume ratio increases. We then developed a logic utilizing the model’s standard deviation (std) specifically for cases exceeding this median. Notably, in these live trading tests, we observed that the model demonstrated predictive power where the increase in volume ratio was higher than the median. This observation offers insights into the practical utility of the model’s uncertainty towards volume spikes in actual market scenarios. To account for sudden jumps in predicted values, we incorporated a constant factor of approximately 0.2, which was multiplied by the standard deviation and added to the predicted values. This constant was empirically adjusted based on observed performance.
@@ 6 Conclusion
In conclusion, our study presents an advancement in the field of intraday volume ratio prediction, integrating transformer architecture with practical trading applications.
Acknowledgments
@@ A.1.1 Data and Features
We utilized execution performance data from September 26, 2023 to November 21, 2023. The features used in the analysis are as follows:
@@ A.1.2 Multiple Regression Analysis Results
The results of the multiple regression analysis are summarized in Table 4.
@@ A.1.3 Interpretation of Results
  • min_bar_low2high_std (x4): The algorithm performs better in conditions of high volatility within minute bars. This implies that IVE is effective in environments with frequent extreme price movements.